OpenStudy (anonymous):

Find the roots of the polynomial equation. x2 – 8x – 48 = 0 A. –12, 4 B. –4, 12 C. –6, 8 D. –8, 6 B?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

\[\huge x^2 – 8x – 48 = 0 \rightarrow x^2 – 12x +4x– 48 = 0\] \[\huge \rightarrow x(x – 12) +4(x– 12) = 0\] \[\huge \rightarrow (x – 12) (x+4) = 0\] i.e. x-12= 0 or x+4 = 0 i.e. x=12 or x=-4 hence x= -4, 12 are the required roots ofthe given quadratic equation. thus option B. –4, 12 is the correct answer. @blueweek77

OpenStudy (anonymous):

Thanks @dpasingh :D